luvvdani1516:

A cylinder has a volume of 300 cubic inches and its height is equal to its radius. The cylinder and a sphere have the same radius. What is the volume of the sphere? FORMULAS: Cylinder- 3.14 X r^2 X h Sphere- 4/3 X 3.14 X r^3

darkknight:

cylinder volume = pi * r^2 * h plug and chug, \[300=\pi*r^2*h\] where \[r=h\] so \(300=\pi*r^2*r\) or \(300=\pi*r^3\) so \(300/ \pi =r^3\) and \[r=\sqrt[3]{300/\pi}\]

darkknight:

sphere volume \[V=4/3*\pi*r^3\] we found r to be \[r=\sqrt[3]{300/\pi}\] plug and chug \[V=4/3*\pi*\sqrt[3]{300/\pi}\] solve for the volume

darkknight:

@rival You don't know what you are doing, I do, don't waste any more brain cells over here

luvvdani1516:

My answer choices are: A) 1200 cubic inches B) 400 cubic inches C) 200 cubic inches D) 100 cubic inches And I got the answer 122.799205. So, which answer choice do I pick?

darkknight:

@darkknight wrote:
sphere volume \[V=4/3*\pi*r^3\] we found r to be \[r=\sqrt[3]{300/\pi}\] plug and chug \[V=4/3*\pi*\sqrt[3]{300/\pi}\] solve for the volume
bruh im stupid, I meant we have to cube r, my bad \[V=4/3*\pi*(\sqrt[3]{300/\pi})^3\]

darkknight:

@luvvdani1516 wrote:
My answer choices are: A) 1200 cubic inches B) 400 cubic inches C) 200 cubic inches D) 100 cubic inches And I got the answer 122.799205. So, which answer choice do I pick?
follow all the steps above, I did not get 122 or anywhere near that

luvvdani1516:

@darkknight wrote:
@luvvdani1516 wrote:
My answer choices are: A) 1200 cubic inches B) 400 cubic inches C) 200 cubic inches D) 100 cubic inches And I got the answer 122.799205. So, which answer choice do I pick?
follow all the steps above, I did not get 122 or anywhere near that
I did and that's what I got.

darkknight:

\[V=4/3*\pi*(\sqrt[3]{300/\pi})^3\] you plugged this in and got 122?

darkknight:

@luvvdani1516 still there?

luvvdani1516:

@darkknight wrote:
\[V=4/3*\pi*(\sqrt[3]{300/\pi})^3\] you plugged this in and got 122?
I keep getting 122. What am I doing wrong? I'm putting everything there into my calculator and I'm still getting the same answer.

darkknight:

darkknight:

or \[4/3*\pi*300/\pi=4/3*300=400\] because the cube root and cube cancels

luvvdani1516:

@darkknight wrote:
I don't understand why I was not getting that answer. But, thanks for the help.

darkknight:

np

surjithayer:

\[r^3=\frac{ 300 }{\pi }\] volume of sphere \[=\frac{ 4 }{ 3 } \pi r^3\] \[=\frac{ 4 }{ 3 }\pi (\frac{ 300 }{ \pi })\] =400 cubic inches.

luvvdani1516:

@surjithayer wrote:
\[r^3=\frac{ 300 }{\pi }\] volume of sphere \[=\frac{ 4 }{ 3 } \pi r^3\] \[=\frac{ 4 }{ 3 }\pi (\frac{ 300 }{ \pi })\] =400 cubic inches.
That's so much easier to understand lol

luvvdani1516:

Thank you for the help everyone!